# Reliability, and Global Privacy

### v3

@@ VIII Convergence
Assumption VIII.2 (Unbiasedness and Bounded Variance).
@@
Assumption VIII.3 (Bounded Dissimilarity).
@@
This combination of assumptions VIII.1VIII.3 is the most relaxed than those commonly adopted in the literature [15] and allows for broad applicability in real-world settings.
Lemma VIII.1 (Accumulation of Local Updates).
With arbitrary local solvers, the accumulated local update on each client after $R_{t_{s}}$ training rounds, each containing $I$ consecutive iterations, is given by
@@
$\sum_{i_{r}=1}^{R_{t_{s}}I}a_{i_{r}}\mathbb{E}\left[\left\lVert\Delta\bm{\Theta}_{k}^{t,R_{t_{s}}I}-\bm{\Theta}^{t-1}\right\rVert^{2}\right]$
@@
where $a_{i_{r}}=a_{i_{r}\bmod I},i_{r}=1,\cdots,R_{t_{s}}I$, and $a_{i_{r}\bmod I}$ corresponds to entries in $\bm{a}$, defined in (4).
@@
By replacing II with $R_{t_{s}}I$ and substituting (57) into [15, Lemma 3], we derive Lemma VIII.1. ∎
Theorem VIII.1.
Given an arbitrary local solver $\bm{a}$, for any algorithms that result in the binary recovery of the global model with failure probability $P_{O}$, if the learning rate is set as $\eta=\frac{1}{G}\sqrt{\frac{M}{T}}$ and the threshold of total training rounds $T$ is large enough, under Assumptions VIII.1VIII.3, the following convergence bound yields with probability $99.74\%$,
$\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t-1})\rVert^{2}\right]\leq\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t-1})\rVert^{2}\right]$
$\leq\frac{2G(\mathcal{L}(\bm{\Theta}^{T})-\mathcal{L}^{\star})}{\lVert\bm{a}\rVert_{1}(1-P_{O})\sqrt{KT}}+\frac{2\lVert\bm{a}\rVert_{1}(1+P_{O})}{\left(1-P_{O}\right)^{2}}\sqrt{\frac{K}{T}}(\sigma^{2}+2\kappa^{2}).$ (59)
where $T_{s}=\min\{n:\sum_{t=1}^{n}R_{t_{s}}\geq T\}$ denotes the total number of successful global update.
@@
Proofs are provided in [32, Appendix G]. ∎
Remark VIII.1 (Role of $P_{O}$).
A large value of $P_{O}$ not only reduces the participation frequency, thereby slowing the convergence rate by $1-P_{O}$, but also introduces a slowdown by $(1-P_{O})^{2}$, due to data heterogeneity and local data variance.
Remark VIII.2 (Role of Privacy).
The convergence analysis demonstrates that the convergence behavior of SecCoGC is independent of its privacy performance, offering a theoretical optimality guarantee under arbitrarily strong privacy levels.
IX Simulations
IX-A Experimental Setups
@@
    Fair-SecCoGC, as described in Section III and IV.
@@
    Private standard FL under unreliable communication, where the Gaussian mechanism is employed.
@@
    Private diversity network codes (DNC)-based FL[24], where the Gaussian mechanism is employed.
The update rule in partial participation in methods (iii), (iv), (v) is non-blind and follows [15, (17)]. That is, the effective step size is not reduced by missing updates.
The MNIST [34] is an easy-level benchmark dataset with clean and standardized format. The CINIC-10 [35], formed by both CIFAR-10 and downsampled ImageNet, is more difficult to train due to greater diversity and real-world noise. To simulate varying levels of data heterogeneity, local datasets are assigned using the Dirichlet distribution. The concentration parameter $\Gamma$ controls the level of heterogeneity among clients, where smaller values of $\Gamma$ indicate a higher imbalance.
@@ IX-B Results and Discussions
Fast Convergence Under Privacy Preservation
When $\lambda$ is set to $0.05$ for MNIST and $0.03$ for CINIC-10, i.e., moderate levels of privacy noise, the private benchmark methods (including both private standard FL and private DNC) can work, but with significantly degraded performance. In addition to slower and less stable convergence, the final accuracy is also substantially reduced. This is because noise increases the amplitude of oscillations during optimization, and the larger step size prevents the final convergence point from accurately approaching the true global optimum. For this reason, even though DNC-based FL nearly ensures full recovery of all masked local models, masking still causes a considerable performance gap between ideal FL with perfect connectivity and DNC-based FL. As for the private standard FL, the perturbation induced by privacy-preserving noise, compounded by unreliable communication, disrupts the optimization trajectory of the masked global model. This results in a more pronounced divergence from the true global optimum, thereby leading to inferior performance compared to private DNC. With an increased privacy level ($\lambda=0.1$ for MNIST and $\lambda=0.05$ for CINIC-10), the utility degradation introduced by privacy noise becomes unacceptable. In this high-privacy regime, private standard FL consistently diverges across all configurations. Despite private DNC’s near-complete removal of communication unreliability, the global model’s accuracy remains significantly degraded, highlighting the intrinsic trade-off between strong privacy guarantees and model utility. In contrast, our proposed SecCoGC consistently approaches ideal performance across all privacy levels. This is because SecCoGC reconstructs the global model exactly whenever recovery is possible. Nonetheless, we also observe slight fluctuations of SecCoGC compared to ideal FL, which is due to the increased variance resulting from the accumulation of local model updates, implied by Lemma VIII.1.
Straggler Mitigation Against Unreliable Communication
@@
X Conclusion
In this work, we introduce SecCoGC, a real-field secure aggregation scheme designed to enhance privacy and reliability under communication uncertainties. Building upon this, we propose Fair-SecCoGC, an extension that ensures fairness in user privacy protection while achieving optimal privacy preservation given a total power constraint of the privacy noise. We formalize the underlying problem, present key construction methods, and analyze their integration within FL frameworks. By ensuring the coordination structure of secret keys and the exact reconstruction of the global model, the proposed schemes support fast convergence and optimal performance under frequent communication disruptions, without requiring prior knowledge of the network topology. Through rigorous theoretical analysis and extensive empirical evaluation, we demonstrate that both SecCoGC and Fair-SecCoGC offer strong robustness to unreliable communication while providing arbitrarily strong privacy guarantees.
@@ X Conclusion
Appendix A Proof of Theorem VI.1
The model of peer-to-peer LDP from client $k$ to client $m$ is similar to the standard Gaussian mechanism with noise $\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$, but with imperfect connectivity $\tau_{m,k}$. The client $m$ should not confidently learn whether the local model update $\Delta\bm{\Theta}_{k}^{t}$ or $\Delta\bm{\Theta}_{k}^{t^{\prime}}$ is contained in the received masked message $\bm{Y}_{k}^{t}$. When client $m$ fails to decode message from client $k$, i.e., $\tau_{m,k}=0$, we have $\bm{Y}_{k}^{t}=0$ regardless of the values of $\Delta\bm{\Theta}_{k}^{t},\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathbb{B}_{D}(R)$, it yields perfect privacy, i.e.,
@@
$\epsilon_{m,k}^{(1)}=\frac{2R}{\lambda}\left[2\log\left(\frac{1.25}{\delta_{m,k}^{(1)}}\right)\right]^{\frac{1}{2}}.$ (62)
Following similar steps in [7, (73)], we have
@@
Appendix B Proof of Theorem VI.2
Now let us proceed to LDP at the relaying client $k$. The relaying client $k$ successfully collects the masked local model updates from clients in $\mathcal{U}_{k}^{t}$ (including itself), computes their partial sum $\bm{W}_{k}^{t}$, and transmits $\bm{W}_{k}^{t}$ to the server. During the transmission, an eavesdropper should not be able to distinguish the participation of client $j$ by observing $\bm{W}_{k}^{t}$.
If client $j\notin\mathcal{U}_{k}^{t}$, then $\bm{W}_{k}^{t,-j}=\bm{W}_{k}^{t}$. So for any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S})\hskip-1.42262pt=$ $\mathrm{P_{r}}(j\in\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})$
$+\mathrm{P_{r}}(j\notin\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\hskip-2.84526pt\mathcal{S}|j\notin\mathcal{U}_{k}^{t}).$ (64)
Consider the network realizations reflected by $\mathcal{U}_{k}^{t}$, such that corresponding variance $\nu_{k}^{t,-j}$ lies within the event $\mathcal{E}\triangleq\{\lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{1}\}$, where $r$ is chosen such that $\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})<\delta^{\prime}$, $\delta^{\prime}\in(0,1)$, then
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})=$ $\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})$
$\hskip-34.1433pt+\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E}^{\mathrm{c}})$
$\hskip-42.67912pt\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E}).$ (65)
@@
$\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})=\hskip-19.91692pt\sum_{\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{1}\end{subarray}}\hskip-19.91692pt\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\mathrm{P_{r}}(\mathcal{A}).$ (66)
@@
$\log\left(\frac{\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}=\bm{z}|i\in\mathcal{A})}{\mathrm{P_{r}}(\bm{W}_{k}^{t}=\bm{z}|i\in\mathcal{A})}\right)$
@@
$\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\leq e^{\epsilon_{k,j}^{\mathcal{A},2}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)},$ (71)
@@
$\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})$ (72a)
$\leq\hskip-14.22636pt\sum_{\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\mathcal{A})\left(e^{\frac{c_{1}}{\sqrt{\nu_{k}^{t,-j}}}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}\right)$ (72b)
$\leq e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\hskip-14.22636pt\sum_{\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\mathcal{A})\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}$ (72c)
$\leq e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)},$ (72d)
@@
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})$
$\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\left(e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}\right)$
$\leq\delta^{\prime}+e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}.$ (73)

### v4

@@ VI Convergence
Assumption VI.2 (Unbiasedness and Bounded Variance).
@@
Assumption VI.3 (Bounded Dissimilarity).
@@
A distinct feature of this convergence analysis is that both the number of consecutive training rounds between two adjacent successful recoveries of the global model and the total number of training rounds are random variables. Moreover, the latter is explicitly dependent on the realizations of the former. Let $t_{s}$ index the $t_{s}$-th successful global model construction, let $R_{t_{s}}$ denote the number of consecutive training rounds the FL system performs before achieving the $t_{s}$-th successful global model construction, and let $T_{s}$ be the total number of successful global updates under the stopping rule, their relationship is drawn in Fig. 2. Given the failure probability of the global model construction $P_{O}$  [^13] A detailed analysis of $P_{O}$ in SecCoGC is identical to the derivations in [35], to which the reader is referred for full detail., the following lemma characterizes their joint behavior.
Figure 2: Illustration of the number of training rounds $T$, the number of consecutive training rounds between two successful global recoveries $R_{t_{s}}$, and the total number of successful global updates $T_{s}$.
Lemma VI.1.
Define $\mathcal{C}_{1}=\left\{T_{s}>\frac{T}{e_{R}}-3\sqrt{\frac{\nu_{R}T}{e_{R}^{3}}}\right\}$, and $\mathcal{C}_{2}\triangleq\left\{\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}<e_{R^{2}}+3\sqrt{\frac{\nu_{R^{2}}}{T_{s}}}\right\}$, we have
Proof.
The proof is provided in Appendix I. ∎
Another lemma concerning the accumulation of local updates for an arbitrary local solver is stated below.
Lemma VI.2.
With an arbitrary choice of local solver, the accumulated local update on each client after $R_{t_{s}}$ consecutive training rounds, each containing $I$ consecutive iterations, is
@@
$\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\mathbb{E}\left[\left\lVert\Delta\bm{\Theta}_{k}^{t,R_{t_{s}}I}-\bm{\Theta}^{t-1}\right\rVert^{2}\right]$
@@
where $a_{i_{s}}=a_{i_{s}\bmod I},i_{s}=1,\cdots,R_{t_{s}}I$, and $a_{i_{s}\bmod I}$ corresponds to entries in $\bm{a}$, defined in (4).
@@
By replacing II with $R_{t_{s}}I$ and substituting (42) into [21, Lemma 3], Lemma VI.2 is derived. ∎
Theorem VI.1.
Given an arbitrary local solver $\bm{a}$, for any algorithm that results in a binary recovery of the global model with failure probability $P_{O}$, if the learning rate is set as $\eta=\frac{1}{G}\sqrt{\frac{M}{T}}$ and the threshold of total training rounds $T$ is large enough, under Assumptions VI.1VI.3, the following convergence bound yields with probability $99.74\%$,
$\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]\leq\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
$\leq\frac{2G(\mathcal{L}(\bm{\Theta}^{0})-\mathcal{L}^{\star})}{\lVert\bm{a}\rVert_{1}(1-P_{O})\sqrt{KT}}+\frac{2\lVert\bm{a}\rVert_{1}(1+P_{O})}{\left(1-P_{O}\right)^{2}}\sqrt{\frac{K}{T}}(\sigma^{2}+2\kappa^{2}).$ (44)
where $T_{s}=\min\{n:\sum_{t_{s}=1}^{n}R_{t_{s}}\geq T\}$ denotes the total number of successful global update.
@@
Proofs are provided in [32, Appendix H]. ∎
Remark VI.1 (Role of $P_{O}$).
A large $P_{O}$ not only reduces the update frequency and slows the convergence rate by a factor of $1-P_{O}$, but also introduces an additional slowdown by $(1-P_{O})^{2}$ due to data heterogeneity and local data variance.
Remark VI.2 (Role of Privacy).
The convergence behavior of SecCoGC is independent of the privacy parameters, ensuring the optimality regardless of privacy levels.
VII Simulations
VII-A Experimental Setups
@@
    SecCoGC, as described in Section III.
@@
    Private standard FL under unreliable communication, which employs the Gaussian mechanism.
@@
    Private diversity network codes (DNC)-based FL[29], which employs the Gaussian mechanism and enables full client-to-client communication among $K$ clients.
The update rule under partial participation (iii), (iv), (v) is set non-blind and follows [29, (17)]. In such a design, the effective step size is not reduced by the number of missing updates.
The MNIST [36] is an easy-level benchmark dataset with a clean and standardized format. The CINIC-10 [37], formed by both CIFAR-10 and downsampled ImageNet, is more difficult to train due to greater diversity and real-world noise. To simulate varying levels of data heterogeneity, local datasets are assigned using the Dirichlet distribution. The concentration parameter $\Gamma$ controls the level of heterogeneity among clients, where smaller values of $\Gamma$ indicate a higher imbalance.
@@ VII-B Results and Discussions
VII-B1 Fast Convergence Under Privacy Preservation
When $\lambda$ is set to $0.05$ for MNIST and $0.03$ for CINIC-10, i.e., moderate levels of privacy noise, the private benchmark methods (including both private standard FL and private DNC) can work, but with significantly degraded performance. In addition to slower and less stable convergence, the final accuracy is also substantially reduced. This is because noise increases the amplitude of oscillations during optimization, and the larger step size prevents the final convergence point from accurately approaching the true global optimum. For this reason, even though DNC-based FL nearly ensures full recovery of all masked local models, masking still causes a considerable performance gap between ideal FL with perfect connectivity and DNC-based FL. As for the private standard FL, the perturbation induced by privacy-preserving noise, compounded by unreliable communication, disrupts the optimization trajectory of the masked global model. This results in a more pronounced divergence from the true global optimum, thereby leading to inferior performance compared to private DNC. With an increased privacy level ($\lambda=0.1$ for MNIST and $\lambda=0.05$ for CINIC-10), the utility degradation introduced by privacy noise becomes unacceptable. In this high-privacy regime, private standard FL consistently diverges across all configurations. Despite private DNC’s near-complete removal of communication unreliability, the global model’s accuracy remains significantly degraded, highlighting the intrinsic trade-off between strong privacy guarantees and model utility. In contrast, our proposed SecCoGC consistently approaches ideal performance across all privacy levels. This is because SecCoGC reconstructs the global model exactly whenever recovery is possible. Nonetheless, we also observe slight fluctuations of SecCoGC compared to ideal FL, which is due to the increased variance resulting from the accumulation of local model updates, implied by Lemma VI.2.
VII-B2 Straggler Mitigation Against Unreliable Communication
@@
VIII Conclusion
In this work, we introduce SecCoGC, a coding-enforced robust secure aggregation method. SecCoGC improves the reliability of zero-sum noise cancellation under unreliable communication, supports arbitrary privacy levels, and preserves the utility of the resulting global model. To conclude, a complete problem formulation of the real-field zero-sum noises is provided, along with the construction conditions and methods. A thorough performance analysis of the proposed SecCoGC in terms of privacy and convergence is provided. Finally, the superiority of the proposed method is validated through numerical comparisons with state-of-the-art benchmarks over multiple datasets.
@@ VIII Conclusion
Due to the page limit, all supporting proofs are detailed in the supplementary file [32], which can be found at https://arxiv.org/abs/2507.07565.
In particular, the appendices are structured as follows.
  • Proofs of Theorem V.1 are in [32, Appendix A].
  • Proofs of Theorem V.2 are in [32, Appendix B].
  • Proofs of Theorem V.3 are in [32, Appendix C].
  • Proofs of Theorem V.4 are in [32, Appendix D].
  • Proofs of Theorem V.5 are in [32, Appendix E].
  • Choices of Bernstein Parameters $r_{1},r_{2},r_{3},r_{4}$ are discussed in [32, Appendix F and G].
  • Proofs of Theorem VI.1 are in [32, Appendix H and I].
Appendix A Proof of Theorem V.1
Definition A.1 ($(\epsilon,\delta)$-LDP).
Definition A.2 (Guassian Mechanism[3]).
Despite the standard Gaussian mechanism $\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$ is applied to the peer-to-peer transmission from client $k$ to client $m$, the correlation between the secret key $\bm{N}_{m}^{t}$ at client $m$ and $\bm{N}_{k}^{t}$ can reveal some information about the unknown secret key. This correlation will weaken privacy to some extent. Lemma IV.4 quantitatively shows the uncertainty reduction reflected by a variance decrease from $\lambda^{2}$ to $\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m}$. By utilizing the correlation between $\bm{N}_{m}^{t}$ and $\bm{N}_{k}^{t}$, the transmission is equivalent to applying the Gaussian mechanism $\mathcal{N}\!\left(\frac{\bm{n}_{m}}{\lambda^{2}}\bm{R}_{m,k},\;(\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m})\cdot\bm{I}_{D}\right)$. The expected value $\frac{\bm{n}_{m}}{\lambda^{2}}\bm{R}_{m,k}$ does not affect the privacy loss, hence it is further equivalent to applying the Gaussian mechanism $\mathcal{N}\!\left(\mathbf{0},\;(\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m})\cdot\bm{I}_{D}\right)$.
When client $m$ fails to decode message from client $k$, i.e., $\tau_{m,k}=0$, we have $\bm{Y}_{k}^{t}=0$ regardless of the values of $\Delta\bm{\Theta}_{k}^{t},\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathbb{B}_{D}(R)$, it yields perfect privacy, i.e.,
@@
$\epsilon_{m,k}^{(1)}=\frac{2R}{\sqrt{\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m}}}\left[2\log\left(\frac{1.25}{\delta_{m,k}^{(1)}}\right)\right]^{\frac{1}{2}}.$ (50)
Following similar steps in [4, (73)], we have
@@
Appendix B Proof of Theorem V.2
Now let us proceed to LDP at the relaying client $k$. The relaying client $k$ successfully collects the masked local model updates from clients in $\mathcal{U}_{k}^{t}$ (including itself), computes their partial sum $\bm{S}_{k}^{t}$, and transmits $\bm{S}_{k}^{t}$ to the server. During the transmission, an eavesdropper should not be able to distinguish the participation of client $j$ by observing $\bm{S}_{k}^{t}$.
If client $j\notin\mathcal{U}_{k}^{t}$, then $\bm{S}_{k}^{t,-j}=\bm{S}_{k}^{t}$. So for any measurable set $\mathcal{S}$, it holds that
$\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S})\hskip-1.42262pt=$ $\mathrm{P_{r}}(j\in\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})$
$+\mathrm{P_{r}}(j\notin\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\hskip-2.84526pt\mathcal{S}|j\notin\mathcal{U}_{k}^{t}).$ (52)
Consider the network realizations reflected by $\mathcal{U}_{k}^{t}$, such that corresponding variance $\nu_{k}^{t,-j}$ lies within the event $\mathcal{E}\triangleq\{\lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{1}\}$, where $r_{1}$ is chosen such that $\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})<\delta^{\prime}$, $\delta^{\prime}\in(0,1)$, then
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})=$ $\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})$
$\hskip-34.1433pt+\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E}^{\mathrm{c}})$
$\hskip-42.67912pt\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E}).$ (53)
@@
$\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})=\hskip-19.91692pt\sum_{\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{1}\end{subarray}}\hskip-19.91692pt\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\mathrm{P_{r}}(\mathcal{A}).$ (54)
@@
$\log\left(\frac{\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}=\bm{z}|i\in\mathcal{A})}{\mathrm{P_{r}}(\bm{S}_{k}^{t}=\bm{z}|i\in\mathcal{A})}\right)$
@@
$\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\leq e^{\epsilon_{k,j}^{\mathcal{A},2}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)},$ (59)
@@
$\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})$ (60a)
$\leq\hskip-14.22636pt\sum_{\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\mathcal{A})\left(e^{\frac{c_{1}}{\sqrt{\nu_{k}^{t,-j}}}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}\right)$ (60b)
$\leq e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\hskip-14.22636pt\sum_{\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\mathcal{A})\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|i\in\mathcal{A})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}$ (60c)
$\leq e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)},$ (60d)
@@
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})$
$\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\left(e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t},\mathcal{E})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}\right)$
$\leq\delta^{\prime}+e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\cdot\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\in\mathcal{U}_{k}^{t})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(2)}.$ (61)